Prove or disprove: For all x ∈Z,if x mod11=5,then x 2 mod11=3.


1
Expert's answer
2022-03-18T01:10:37-0400

Solution


Given that


"x" mod11 "= 5"


Then, for some value of "k", we can write x mod11= 5, as


"x = 11k + 5"


Then


"x^{2}=(11k+5)(11k+5)"


"x^{2}=121k^2+55k+55k+25"


"x^{2}=121k^2+110k+25"


Now,


"x^2" mod 11"=(121k^2+110k+25)" mod 11


"x^2" mod 11"=121 k^2" mod 11 + "110k" mod 11+ "25" mod 11


"x^2" mod 11"=0" + 0+ 3 because "25=2\times11+3"


Therefore,


"x^2" mod 11 "= 3"